OpenStudy (anonymous):

What is the simplified form of (10/(8xy) multiplied by (4x^2)/(3y)? Can anybody solve this?

OpenStudy (anonymous):

can you help me @satellite73

OpenStudy (anonymous):

\[\frac{ 10 }{ 8xy }\times \frac{ 4x ^{2} }{ 3y }\] like this?

OpenStudy (anonymous):

is that the correct equation?

OpenStudy (anonymous):

Yes except the 10 is actually supposed to be a 1 sorry about the typo lol.

OpenStudy (anonymous):

just do numerator times numerator and denominator times denominator then cancel variable which are the same

OpenStudy (anonymous):

so it would be 2x/3y^2

OpenStudy (anonymous):

\[\frac{ 1 }{ xy }\times \frac{ 4x ^{2} }{ 3y }=\frac{ 4x ^{2} }{ 3xy ^{2} }\]

OpenStudy (anonymous):

underneath the 1 it should be 8xy you only put xy

OpenStudy (anonymous):